OpenStudy (anonymous):

derivative of y = sin(tan2x)

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

First, a quick review. The outermost / most important function here is the sine. Think of the input to this sine functin as u, a separate function of x, which in this case happens to be tan 2x. You with me so far?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

what rule to be apply in this problem?

OpenStudy (mathmale):

1) derivative of the sine function 2) derivative of the tangent function 3) two applications of the chain rule

OpenStudy (anonymous):

oh its hard to solve :3

OpenStudy (mathmale):

Your outermost function is the sine: y = sin ( tan 2x ) ) looks like y = sin u. Your inner function (the argument of the sine) is u = tan 2x.

OpenStudy (anonymous):

what should i do 1st?

OpenStudy (mathmale):

We're going to do some review first. What is the derivative of the sine function? What is the derivative of y = sin x?

OpenStudy (anonymous):

y' = cos x

OpenStudy (mathmale):

Good. Suppose y = tan x. What is the derivative, y ' , of this function? If possible look it up before you answer.

OpenStudy (anonymous):

y' = sec^2 x ?

OpenStudy (mathmale):

Good. Making progress! OK. Look at y = sin (tan x ). (I did not forget the '2x') this looks like y = sin u, where u = tan x, right? Since u is a separate function of x, we must use the Chain rule. First, we differentiate the "sin u" and then we multiply the result by the derivative of u. Clear, or need mor explaination or examples?

OpenStudy (anonymous):

can u give some example?

OpenStudy (mathmale):

Let's start with y = sin ( tan x ). Use these rules: 1) the derivative of the sine function is the cosine function: [ sin x ] ' = cos x and [ sin u ] ' = cos u. 2) If u happens to be a function of x (which is the case in the problem you've posted), then we have to apply the CHAIN RULE and differentiate u as well, with respect to x: [ sin u ] ' = [ cos u ] * u '

OpenStudy (mathmale):

Example: y = sin ( cos x ); the derivative is found by applying the Chain Rule: y ' = cos ( cos x ) * [ cos x ] ' = cos ( cos x ) * ( -sin x ). You can't simplify that further. Leave it as is.

OpenStudy (mathmale):

Let's go back to y = sin ( tan x ). Here u = tan x and u ' = ( sec x )^2. Therefore, y ' = cos ( tan x ) * ( sec x )^2. That's the answer. Questions?

OpenStudy (anonymous):

the u' = sec x ^ 2 or u ' = sec^2 x ? which is true?

OpenStudy (mathmale):

If u = tan x, then u ' = (sec x)^2.

OpenStudy (anonymous):

okay whats next?

OpenStudy (mathmale):

First: decide whether or not you're OK with y ' = cos ( tan x ) * ( sec x )^2. as the derivative of y = sin (tan x). If you are, fine; if not, tell me what needs further explanation.

OpenStudy (anonymous):

i understand why it become y' = cos ( tan x ) * ( sec x )^2

OpenStudy (mathmale):

OK, great. Now we're almost ready to solve the original problem that you've posted: derivative of y = sin(tan2x) What is the derivative of tan 2x ? Hint: Be certain to use the Chain Rule.

OpenStudy (anonymous):

what is the equivalent of tan 2x?

OpenStudy (mathmale):

There's a formula for that, a trig identity, and I'd be happy to help you find it, but right now you don't really need that trig identity. Hint: the derivative of tan x is (sec x)^2. Go one step further and find the derivative of tan 2x using the chain rule.

OpenStudy (anonymous):

tan 2x' = 2sec^2 (2x) ?

OpenStudy (mathmale):

Very close. WRite [tan 2x] ' instead of tan 2x ' for clarity.

OpenStudy (anonymous):

ok ok

OpenStudy (mathmale):

The right side of your equation is entirely appropriate; you have enclosed that '2x' inside parentheses as you should. Where did the 2 in front of [sec(2x}] come from? It is correct, but I'd like to know why you write your answer that way.

OpenStudy (anonymous):

i'm only substitute the value of 2

OpenStudy (mathmale):

Actually, Jay, you are differentiating the argument '2x'; the derivative of that is 2. According to the Chain Rule you must differentiate (2x) and multiply the derivative of tan (2x) by your result (2).

OpenStudy (mathmale):

Please fix your tan 2x' = 2sec^2 (2x) Hint: review our discussion, above.

OpenStudy (anonymous):

[tan 2x] ' = 2sec^2 (2x)

OpenStudy (mathmale):

Great. Perfect. You're done.

OpenStudy (anonymous):

whats next

OpenStudy (anonymous):

can you summarize the flow of solution ?

OpenStudy (mathmale):

Look at what you initially posted: "derivative of y = sin(tan2x)" Then look at what you ended up with: [tan 2x] ' = 2sec^2 (2x) Explain this in words.

OpenStudy (mathmale):

What was the goal of this math prob lem?

OpenStudy (anonymous):

to get the derivative of y

OpenStudy (mathmale):

Right, and you have done that. The only thing missing is a label for y our result:\[\frac{ dy }{ dx }=[\tan 2x] ' = 2\sec^2 (2x) \]

OpenStudy (mathmale):

or you could have used y ' as your label. In either case you are done with this differentiation and your result is correct.

OpenStudy (anonymous):

what could be happen to sin?

OpenStudy (mathmale):

Review: "derivative of y = sin(tan2x)" 1. differentiate sin (tan 2x) with respect to tan 2x: you'll get cos (tan 2x). 2. differentiate that tan 2x; you'll get 2(sec 2x)^2. 3. multiply together the results of (1) and (2), above. Answer: y ' = cos (tan 2x) * 2(sec 2x)^2.

OpenStudy (mathmale):

It's essential that you be able to do this yourself, without help. You can ask all the questions you want. Review our discussion. List, review and apply the rules of differentiation I've shared with you: Product rule Chain rule Implicit differentiation Power rule Power rule with chain rule Constant coefficient rule Derivative of a constant and so on. It's imperative you have a list of these rules and that you can recognize when to apply each of them.

OpenStudy (anonymous):

thanks man

OpenStudy (mathmale):

My pleasure. Altho I need to get off OpenStudy now, i'd be delighted to work with you again. You've written down my e-mail address? I have your yahoo address.

OpenStudy (mathmale):

Are you located in a country other than the USA? If so, which one?

OpenStudy (anonymous):

i'm from philippines.

OpenStudy (mathmale):

so your first language is Tagalog?

OpenStudy (anonymous):

i can't add you in facebook

OpenStudy (anonymous):

yeah. that's why im not good in english

OpenStudy (mathmale):

You can always improve your English. Let's worry about the Facebook thing later on, ok? for now we can communicate through OpenStudy or e-mail. \